OpenStudy (anonymous):

A triangular prism has a volume of 5 m3. If each dimension of the prism is tripled, what is the new volume of the prism? Answers 135 m3 45 m3 15 m3 405 m3

Directrix (directrix):

135 m3 The cube of the scale factor (1 to 3) of the two prisms equals the ratio of their volumes. 1/27 = 5/V V = 135